OpenStudy (anonymous):

@SolomonZelman Find the indicated real nth root(s) of a. Thank you for any help! 1. n=4, a=0 2. n=7, a=128 3. n=5, a=0

OpenStudy (anonymous):

real nth root... just find all the nth roots and throw out all the imaginary ones

OpenStudy (anonymous):

Okay, can you show me how? =)

OpenStudy (anonymous):

what math class are you in?

OpenStudy (anonymous):

I'm in algebra

OpenStudy (anonymous):

are you in a unit about imaginary numbers?

OpenStudy (anonymous):

No, my teacher gives us problems that we aren't learning in class.

OpenStudy (anonymous):

have you learned about imaginary numbers before?

OpenStudy (anonymous):

(just wondering what depth I need to explain it to)

OpenStudy (anonymous):

I have but I can't remember how to find the real nth root.

OpenStudy (anonymous):

I remember most of it so just the nth root!

OpenStudy (anonymous):

for the real nth root yes

OpenStudy (anonymous):

so x^n=a

OpenStudy (anonymous):

Okay, thank you so much!

OpenStudy (anonymous):

do you want me to teach you or remind you how to find complex nth roots?

OpenStudy (anonymous):

Can you remind me?

OpenStudy (anonymous):

So, for the first problem would the answer be x^4=0?

OpenStudy (anonymous):

the answer would be whatever x would be if x^4=0

OpenStudy (anonymous):

complex nth roots are found using:\[\sqrt[n]{r(\cos (\theta) + i \sin (\theta)}=\sqrt[n]{r}(\cos(\frac{ \theta ±2\pi }{ n })+ i \sin (\frac{ \theta ±2\pi }{n})\]

OpenStudy (anonymous):

have you ever seen something like that before?

OpenStudy (anonymous):

if not, you probably don't have to know it

OpenStudy (anonymous):

I haven't seen that before

OpenStudy (anonymous):

=(

OpenStudy (anonymous):

ok, you don't have to know it

OpenStudy (anonymous):

I learned that in pre-calculus, so...

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

you're fine

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

just forget that

OpenStudy (anonymous):

Well, thanks for your help